1,000 (3,500) + 200 (0) + 1,900 (85% P) + 100 (P) = 11,100,000

Answer by checkley71(8403)

1,000*3,500+200*0+1,900*.85P+100P=11,100,000
3,500,000+0+1,615P+100P=11,100,000
1,715P=11,100,000-3,500,000
1,715P=7,600,000
P=7,600,000/1,715
P=4,431.4868 ANSWER.